40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 18 18:16:26 2014, Nichin, Update, Electronics, Changes to the PD frequency response measurement system  
    Reply  Wed Jun 18 21:53:48 2014, Koji, Update, Electronics, Changes to the PD frequency response measurement system  
Message ID: 10062     Entry time: Wed Jun 18 18:16:26 2014     Reply to this: 10065
Author: Nichin 
Type: Update 
Category: Electronics 
Subject: Changes to the PD frequency response measurement system  

[Nichin, Eric G, Koji]

Continuing out work from elog:10037, we wanted to check if the frequency response of AS55. Having figured out exactly how to use the Laser diode controller (LDC 3744C), we hooked up a fiber power meter to the optical fiber illuminating AS55 (that we disconnected from its mount last Friday ) and raised up the current to 150mA to get almost 0.8mW power reading.

When aligning the fiber to illuminate the PD, we found that the beam was pretty wide. So we pulled out the collimator and tweaked it to get a focused beam. The fiber was mounted back and was aligned to get a maximum DC reading. The multimeter readout 30mV finally. Taking the transimpedence as 200ohm approx., the hot current is about 1.5mA.

Network analyzer was now connected to the modulation input of the laser and the RF output from REF DET and AS55 (inputs to RF switch at rack 1Y1) were connected as input to measure the transfer function. We got just noise on the scope of NA. So, then we tried REFL33 as the Input and still got nothing (We were also not sure if this PD was properly illuminated, we did not check). However the REF DET was giving a nice response on the scope. Turns out all the PDs were disconnected form the Demodulator (D990511) on rack 1Y2.

On closer inspection the RF cable between domodulator and RF switch that was labelled AS55 had a loose SMA connector at the switch end. I will have to fix that tomorrow . For the time being Koji connected the cable labelled REFL33 to the AS55 demodulator and we finally got a response form the AS55 PD on the NA. However no readings were recorded. The power supply to REF DET was turned off in the end as Eric G claimed that it has been ON for almost a year now, which is not a good thing. Also, we removed the modulation input from NA to the diode laser and terminated the input with a 50ohm terminator.

We planned to pull out and check each and every RF cable (especially the SMA ends for faulty soldering and loose connections) and fix/ replace them as needed.

ELOG V3.1.3-